# Pseudo-Bifurcations in Stochastic Non-Normal Systems

### v2

@@ Abstract
By measuring the lead-lag correlation between the system’s projection along $\hat{\mathbf{e}}_{n}$ and the mean-field state $\langle\mathbf{x}\rangle$, we observe that the former leads the later by about $0.1$ seconds (Figure 4, bottom-right). This suggests that epileptic seizures may not only be driven by non-normal dynamics but also that the onset of transient episodes could be anticipated by measuring the system’s exposure to non-normal components. This insight offers a promising pathway for enhancing the early detection and potential control of seizure events, while also paving the way for managing self-induced instabilities in complex systems more broadly [44].
Reexamining Criticality
@@
Naturally, we must apply our framework with caution. It is well-established that many systems experience true bifurcations, where crossing a critical threshold leads to a fundamental change in behavior. Such transitions result in a qualitative shift in dynamics, which is both observable and empirically validated. Beyond the cases of desertification [10] and lake eutrophication [11] previously mentioned, coral reefs present a notable example: under stressors like pollution, overfishing, and climate change, reefs can cross a threshold, transitioning from coral-dominated to algae-dominated states. This shift is often irreversible, as the algae-dominated state hinders coral regrowth through altered ecological dynamics [45]. Similarly, certain fisheries, such as the Newfoundland cod fishery, have experienced abrupt collapses due to overfishing; once population levels dropped below a critical threshold, recovery became impossible, disrupting both ecological and economic systems [46].
@@
However, in many cases, the assumption that a system is nearing a critical point remains largely hypothetical, with empirical evidence still inconclusive. For instance, there is ongoing debate over whether epileptic seizures constitute a critical transition in brain dynamics. While some studies suggest that the brain operates near a critical point, others find the evidence inconclusive [18, 17]. Our findings align with the latter perspective, reinforcing the view that seizures do not necessarily correspond to critical transitions.
@@
Similarly, the Atlantic Meridional Overturning Circulation (AMOC) is suspected of approaching a bifurcation point that could lead to abrupt climate shifts, yet definitive observations remain lacking [19]. It is also possible that warmer interglacial periods observed in Earth’s climate history do not result from crossing a critical bifurcation point but instead reflect non-normal transients within an otherwise glacial climate regime. In these examples, the system’s proximity to a critical point is speculative, and a re-evaluation through the lens of non-normal dynamics could offer new insights.
Conclusions
@@
By focusing on non-normal dynamics, we have proposed an alternative explanation for early-warning signals traditionally attributed to bifurcations in complex systems. We demonstrated that systems with a high degree of non-normality can exhibit transient repulsive dynamics – called pseudo-bifurcations – which give rise to the same statistical signatures as true bifurcations, such as critical slowing down and increased variance. Importantly, these patterns do not necessarily indicate the proximity to a critical bifurcation point. Our empirical analysis of EEG recordings during epileptic seizures further supports this hypothesis, showing that the brain’s non-normal dynamics are prominently dominating epileptic seizures. These findings suggest that many systems previously believed to be approaching criticality may, in fact, be at long-term equilibrium but exhibit strong non-normal transients. This calls for a reexamination of systems previously surmised to operate near criticality and offers a new perspective on the anticipation and control of instabilities in dynamical complex systems.
Acknowledgements
This work is partially supported by the National Natural Science Foundation of China (Grant No. T2350710802 and No. 72350410487), Shenzhen Science and Technology Innovation Commission Project (Grant No. GJHZ202107-05141805017 and No. K23405006), and the Center for Computational Science and Engineering at Southern University of Science and Technology.

### v3

@@ Results
By measuring the lead-lag correlation between the system’s projection along $\hat{\mathbf{e}}_{n}$ and the mean-field state $\langle\mathbf{x}\rangle$, we observe that the former leads the later by about $0.1$ seconds (Figure 4, bottom right). This suggests that epileptic seizures may not only be driven by non-normal dynamics but also that the onset of transient episodes could be anticipated by measuring the system’s exposure to non-normal components. This insight offers a promising pathway to improve the early detection and possible control of seizure events, while also paving the way for the management of self-induced instabilities in complex systems more broadly [44].
Discussion
@@
Naturally, we must apply our framework with caution. It is well-established that many systems experience true bifurcations, where crossing a critical threshold leads to a fundamental change in behavior. Such transitions result in a qualitative shift in dynamics that is both observable and empirically validated. Beyond the previously mentioned cases of desertification [10] and lake eutrophication [11], coral reefs provide a notable example: Under stressors such as pollution, overfishing, and climate change, reefs can cross a threshold, transitioning from coral-dominated to algae-dominated states. This shift is often irreversible, as the algae-dominated state hinders coral regrowth through altered ecological dynamics [45]. Similarly, certain fisheries, such as the Newfoundland cod fishery, have experienced abrupt collapses due to overfishing; once population levels dropped below a critical threshold, recovery became impossible, disrupting both ecological and economic systems [46].
@@
However, in many cases, the assumption that a system is nearing a critical point remains largely hypothetical, with empirical evidence still inconclusive. For example, there is ongoing debate on whether epileptic seizures constitute a critical transition in brain dynamics. Although some studies suggest that the brain operates near a critical point, others find the evidence to be inconclusive [18, 17]. Our findings align with the latter perspective, reinforcing the view that seizures do not necessarily correspond to critical transitions.
@@
Similarly, the Atlantic Meridional Overturning Circulation (AMOC) is suspected of approaching a bifurcation point that could lead to abrupt climate shifts, yet definitive observations remain lacking [19]. It is also possible that warmer interglacial periods observed in Earth’s climate history do not result from crossing a critical bifurcation point but instead reflect non-normal transients within an otherwise glacial climate regime. In these examples, the system’s proximity to a critical point is speculative, and a reevaluation through the lens of non-normal dynamics could offer new insights.
@@
In summary, by focusing on non-normal dynamics, we have proposed an alternative explanation for early-warning signals traditionally attributed to bifurcations in complex systems. We demonstrated that systems with a high degree of non-normality can exhibit transient repulsive dynamics - called pseudo-bifurcations - which give rise to the same statistical signatures as true bifurcations, such as critical slowing down and increased variance. Importantly, these patterns do not necessarily indicate proximity to a critical bifurcation point. Our empirical analysis of EEG recordings during epileptic seizures further supports this hypothesis, showing that the brain’s non-normal dynamics are prominently dominating epileptic seizures. These findings suggest that many systems previously believed to be approaching criticality may, in fact, be at long-term equilibrium but exhibit strong non-normal transients. This calls for a reexamination of systems previously assumed to operate near criticality and offers a new perspective on the anticipation and control of instabilities in dynamical complex systems.
Materials and Methods
Dimensional Reduction of Non-Normal Dynamics
Because matrix $\mathbf{A}$ that governs Eq. (1) is assumed diagonalizable, there exists a matrix $\mathbf{P}$ which is the eigenbasis transformation of $\mathbf{A}$ such that $\mathbf{A}=\mathbf{P}\boldsymbol{\Lambda}\mathbf{P}^{-1}$ where $\boldsymbol{\Lambda}=-\text{diag}\left(\lambda_{1},\cdots,\lambda_{n}\right)$ is the diagonal matrix composed of the eigenvalues of $\mathbf{A}$. Here, we assume that $\lambda_{i}>0~{}\forall i$ and we explicitly set the minus sign to indicate that the eigenvalues are negative. If $\mathbf{A}$ is normal, that is, if $\mathbf{A}\mathbf{A}^{\dagger}=\mathbf{A}^{\dagger}\mathbf{A}$, it would be the case that $\mathbf{P}$ is unitary, that is $\mathbf{A}$ can be diagonalized by means of a rotation. But in general, we assume here that $\mathbf{A}$ is non-normal, and hence $\mathbf{P}$ is not unitary. Quite generally, a matrix $\mathbf{P}$ can be represented as a sequence of a rotation, a rescaling, and another rotation, via the singular value decomposition (SVD) given by $\mathbf{P}=\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{\dagger}$ where $\mathbf{U}$ and $\mathbf{V}$ are unitary matrices, $\mathbf{V}^{\dagger}$ is the Hermitian conjugate of $\mathbf{V}$ and $\boldsymbol{\Sigma}=\text{diag}(\sigma_{1},\cdots,\sigma_{n})$ is a diagonal matrix of the singular values $\{\sigma_{i}\}$. Without loss of generality, we can assume that $\sigma_{1}\geqslant\sigma_{2}\geqslant\ldots\geqslant\sigma_{n}$ because a permutation of a unitary matrix remains unitary. Plugging $\mathbf{P}$ into the above expression for $\mathbf{A}$ yields
$\mathbf{A}=\mathbf{P}\boldsymbol{\Lambda}\mathbf{P}^{-1}=\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{\dagger}\boldsymbol{\Lambda}\mathbf{V}\boldsymbol{\Sigma}^{-1}\mathbf{U}^{\dagger}=\sum_{i,j=1}^{n}\frac{\sigma_{i}}{\sigma_{j}}\lambda_{ij}\mathbf{\hat{u}}_{i}\mathbf{\hat{u}}_{j}^{\dagger}$ (6)
with $\lambda_{ij}=-\sum_{k=1}^{n}\lambda_{k}v_{ik}v_{jk}^{*},$ and where $\mathbf{u}_{i}$ and $\mathbf{v}_{i}$ denote the $i^{\text{th}}$ column of the unitary matrix $\mathbf{U}$ and $\mathbf{V}$, respectively. Further, $v_{ij}$ is the element of $\mathbf{V}$ at the $i^{\text{th}}$ column and $j^{\text{th}}$ row and $v_{ij}^{*}$ is the complex conjugate of $v_{ij}$. Expression (6) shows how $\mathbf{A}$ can be decomposed into $n^{2}$ rank-1 projections $\hat{\mathbf{u}}_{i}\hat{\mathbf{u}}_{j}^{\dagger}$ modulated by a factor $\frac{\sigma_{i}}{\sigma_{j}}\lambda_{ij}$ We now assume that only the last dimension has a singular value significantly smaller than $1$, while all other dimensions are roughly equal to one, that is $1=\sigma_{1}\gtrsim\sigma_{2}\gtrsim\ldots\gtrsim\sigma_{n-1}\gg\sigma_{n}$. This assumption essentially encapsulates the non-normal behavior along one component. We show in the SM Appendix that our results remain qualitatively similar as long as the ratio of the second last to last singular value is large. The condition number of matrix $\mathbf{A}$ is defined as $\kappa=\sigma_{1}/\sigma_{n}$. For $i=2,\ldots,n-1$, it is convenient to write $\sigma_{i}=(1-\delta_{i})\sigma_{1}$ with $\delta_{i}\ll 1$ such that $\sigma_{i}$ is slightly less than but close to $\sigma_{1}=1$. To leading order, this yields
$\frac{\sigma_{i}}{\sigma_{j}}=\begin{cases}1+(\delta_{j}-\delta_{i})+\mathcal{O}(\delta_{i}\delta_{j},\delta_{j}^{2})&\text{if }i,j\neq n\\ \kappa(1-\delta_{i})&\text{if }i\neq n,j=n\\ \kappa^{-1}+\mathcal{O}(\kappa^{-1}\delta_{j})&\text{if }i=n,j\neq n\\ 1&\text{if }i=j=n.\end{cases}$ (7)
Defining further $\delta^{2}=\sum_{i}\delta_{i}^{2}$ and plugging these approximations into (6) allows us to decompose the matrix $\mathbf{A}$ as
$\mathbf{A}=\mathbf{A}_{N}+(\kappa-1)a_{nr}\hat{\mathbf{r}}\hat{\mathbf{n}}^{\dagger}+\kappa\delta\mathbf{e}\hat{\mathbf{n}}^{\dagger}(\kappa^{-1}-1)a_{nr}\hat{\mathbf{n}}\hat{\mathbf{r}}^{\dagger}+\delta\mathbf{E}+\mathcal{O}(\delta^{2},\kappa^{-1}\delta),$ (8)
where $\mathbf{A}_{N}=\mathbf{U}\mathbf{V}^{\dagger}\boldsymbol{\Lambda}\mathbf{V}\mathbf{U}^{\dagger},$ is the “normal component” of $\mathbf{A}$. Moreover, we have defined the non-normal mode $\hat{\mathbf{n}}=\hat{\mathbf{u}}_{n}$, and the reaction mode $\hat{\mathbf{r}}$ via $a_{nr}\hat{\mathbf{r}}=\sum_{i=1}^{n-1}\lambda_{in}\hat{\mathbf{u}}_{i}$ with $a_{nr}=(\sum_{i=1}^{n-1}|\lambda_{in}|^{2})^{1/2}$, both of which are unit vectors. The matrix $\mathbf{E}$ and vector $\mathbf{e}$ are given by $\mathbf{E}=\sum_{i,j=1}^{n-1}(\delta_{j}-\delta_{i})\lambda_{ij}\hat{\mathbf{u}}_{i}\hat{\mathbf{u}}_{i}^{\dagger}$ and $\mathbf{e}=\sum_{i=1}^{n-1}\frac{\delta_{i}}{\delta}\lambda_{in}\hat{\mathbf{u}}_{i}$ and encapsulate the comparatively small interactions resulting from the anisotropy of the small deviation $\delta_{i}$ from non-normality along the $n-1$ dimensions. We can see from equation (8) that $\mathbf{A}$ is decomposed into a normal part as well as off-diagonal parts which are dominated by a term of order $\kappa$. Since these dominating interactions arise along $\hat{\mathbf{r}}$ and $\hat{\mathbf{n}}$, to a good approximation, we can restrict our study to these two dimensions, and project the matrix $\mathbf{A}$ into $(\hat{\mathbf{n}},\hat{\mathbf{r}})$ to obtain system (2). The exact form of $\alpha_{n}$ and $\alpha_{r}$, as well as a derivation that these coefficients are of order one, is found in the SM.
The eigenvalues $\lambda_{\pm}$ of system (2) are $\lambda_{\pm}=-\delta_{+}\mp\sqrt{1+\delta_{-}^{2}}$ where $\delta_{\pm}=\frac{\alpha_{r}\pm\alpha_{n}}{2}$ and it holds $\lambda_{+}>\lambda_{-}>0$. The system is stable if and only if $\lambda_{-}<0$, which implies that $\alpha_{n}\alpha_{r}\geqslant 1$. The associated eigenvectors are given by
$\mathbf{v}_{\pm}=\frac{1}{\sqrt{\mu_{\mp}^{2}+\kappa^{2}}}\begin{pmatrix}\mu_{\mp}\\ \kappa\end{pmatrix}\quad\text{with }\mu_{\pm}=\delta_{-}\pm\sqrt{1+\delta_{-}^{2}}.$ (9)
This shows that, as $\kappa$ and hence the non-normality increases, the two eigenvectors become closer and closer to being collinear. We further detail in the SM that, in our set-up, the matrix condition number $\kappa$ well captures the degree of non-normality that is otherwise typically described via Henrici’s departure from normality.
Non-Normal Statistics
[…]
$\mathbf{C}_{t}=\frac{B^{2}}{4\Delta^{2}}\begin{pmatrix}F_{-}(t)+\kappa^{-2}F(t)&&\kappa G_{-}(t)+\kappa^{-1}G_{+}(t)\\ \kappa G_{-}(t)+\kappa^{-1}G_{+}(t)&&F_{+}(t)+\kappa^{2}F(t)\end{pmatrix}$ (10)
with the functions $F_{\pm}$ and $G_{\pm}$ given in the SM, which are of order one. The variance of the system is dominated by its largest eigenvalue, which we call the effective variance of the system $c_{\text{eff}}(t)$. In the limit of large $\kappa$, it is given by $c_{\text{eff}}(t)=\frac{B^{2}}{4\Delta^{2}}\left[\kappa^{2}F(t)+F_{-}(t)+\frac{G_{-}(t)^{2}}{F(t)}\right]$ which grows like $\kappa^{2}$ to leading order. On the other hand, the second eigenvector is found to be asymptotically independent of $\kappa$ (see SM). In other words, in a highly non-normal system where $\kappa$ is large, most of the dynamics is driven by a single component for which the asymptotic variance grows as $\kappa^{2}$ to leading order. Similarly, we detail in the SM that the system’s lead-lag autocorrelation $C_{r}(\tau)$ asymptotically decays like
$\frac{d}{d\tau}\ln C_{r}(\tau)=-\begin{cases}\frac{\lambda_{+}\lambda_{-}(\lambda_{+}+\lambda_{-})(\lambda_{+}-\lambda_{-})^{2}}{\kappa^{2}(\lambda_{+}-\lambda_{-})^{2}+(\lambda_{+}+\lambda_{-})(\lambda_{-}\mu_{-}^{2}+\lambda_{+}\mu_{+}^{2})+4\lambda_{+}\lambda_{-}}&\tau\rightarrow 0\\ \lambda_{-}&\tau\rightarrow\infty.\end{cases}$ (11)
In the limit of large times, the correlation decays with $\lambda_{-}$ as anticipated. However, the instantaneous rate of decay tends toward zero as $\kappa$ increases, which implies that the lead-lag correlation is near unity for small lags and hence that the system behaves quasi deterministically over short timescales when $\kappa$ is large.
Representation as a Mean-Reversion Process
[…]
$\theta_{t}$ $\propto 2\lambda_{+}(\kappa^{2}+\mu_{-}^{2})e^{-2\lambda_{+}t}+\lambda_{-}(\kappa^{2}+\mu_{+}^{2})e^{-2\lambda_{-}t}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}-(\lambda_{+}+\lambda_{-})(\kappa^{2}-1)e^{-(\lambda_{+}+\lambda_{-})t}.$ (12)
We can now ask whether there exist intermediate times for which $\theta_{t}<0$, that is times where the mean-reversion rate temporarily becomes repulsive. To this end, we look for the roots of (12), which are given by
$\tau_{\pm}=\frac{(\lambda_{+}+\lambda_{-})(\kappa^{2}-1)}{2\lambda_{+}(\kappa^{2}+\mu_{-}^{2})}\pm\frac{1}{2\lambda_{+}(\kappa^{2}+\mu_{-}^{2})}\sqrt{L}$ (13)
with $L=(\lambda_{+}+\lambda_{-})^{2}(\kappa^{2}-1)^{2}-4\lambda_{-}\lambda_{+}(\kappa^{2}+\mu_{+}^{2})(\kappa^{2}+\mu_{-}^{2}).$ If the system is close to normality, $0<\kappa-1\ll 1$, $L$ becomes negative, implying no real solution, and the system remains locally stable with $\theta_{t}>0~{}\forall t$. However, for highly non-normal systems ($\kappa\gg 1$), a real positive solution $\tau_{+}$ exists, showing that there exists a time interval where $\theta_{t}\leqslant 0$.
We can also pinpoint the minimal value $\kappa=\kappa_{c}$ required for $\theta_{t}$ to exhibit transiently negative values. To this end, we determine the value of $\kappa_{c}$ at which $\tau_{+}=\tau_{-}$, that is, the value at which $\theta_{t}$ just touches 00 without crossing below (see the middle plot of Figure 2). It holds that $\tau_{+}=\tau_{-}$ when the square root of (13) is zero, which is the case for
$\kappa_{c}=\sqrt{1+2\lambda_{+}\lambda_{-}+2\sqrt{\lambda_{+}\lambda_{-}(1+\lambda_{+}\lambda_{-})}}.$ (14)
Empirical Estimation of Non-Normal and Reaction Modes
For our application to the brain connectome, we estimate the matrix $\mathbf{A}$ of (1) as $\hat{\mathbf{A}}$ obtained by fitting a VAR(1) process to the EEG data over rolling windows of 40 seconds. In principle, we could then proceed by first determining the transition matrix $\mathbf{P}$ that diagonalizes $\hat{\mathbf{A}}$, and subsequently calculating the SVD $\mathbf{P}=\mathbf{U}\Sigma\mathbf{V}^{\dagger}$ to identify the non-normal mode $\hat{\mathbf{n}}\sim\mathbf{u}_{n}$ and reaction mode $\hat{\mathbf{r}}\sim\sum_{i=1}^{n-1}\mathbf{u}_{i}$. However, as we detail in the SM, this approach is highly sensitive to noise, especially when $\mathbf{A}$ is non-normal. Therefore, we propose an alternative approach that consists in estimating the non-normal and reaction component directly from $\hat{\mathbf{A}}$ as
where $\mathbb{S}^{n-1}$ is the unit sphere in $n$ dimensions. Here, $\langle\mathbf{v},\hat{\mathbf{A}}\mathbf{w}\rangle$ represents the inner product between the vector $\mathbf{v}$ and the result of applying the matrix $\hat{\mathbf{A}}$ to the vector $\mathbf{w}$. Thus, it measures the degree of alignment or projection of the transformed vector $\hat{\mathbf{A}}\mathbf{w}$ on $\mathbf{v}$. Equation (15) directly captures the intuition behind the reduced equation (2) whereby the non-normal component maximally projects onto the reaction component. From (2), we further see that $\kappa$ can be estimated directly from $\hat{\mathbf{r}}$ and $\hat{\mathbf{n}}$ via $\kappa=\sqrt{\left|\langle\hat{\mathbf{r}},\hat{\mathbf{A}}\hat{\mathbf{n}}\rangle\right/\left|\langle\hat{\mathbf{n}},\hat{\mathbf{A}}\hat{\mathbf{r}}\rangle\right|}$ thus allowing us to completely bypass the need for estimating the singular values which are numerically unstable. In the SM, we test our empirical procedure on synthetic data, finding that our procedure faithfully recovers the non-normal and reaction mode, and estimates $\kappa$ more precisely than when calculating it directly via singular value decomposition.